# RoboCar: A Rapidly Deployable Open-Source Platform for Autonomous Driving Research

### v1

@@ I Introduction
This research introduces RoboCar, an open-source research platform developed at the University of Luxembourg. RoboCar provides a modular, cost-effective autonomous driving solution tailored for, but not limited to, the 2018 KIA Soul EV (see Figure 1), enabling advanced research and practical applications in autonomous driving technologies.
@@ II Related work
Finally, the European project PRYSTINE[12] used the same vehicle platform as RoboCar to enhance automotive architectures for automated driving by focusing on fail-operational systems. As the main focus of this project was to investigate fail-operational perception based on sensor fusion, it did not implement a full autonomous driving stack, nor has the implementation been open-sourced.
@@
The research outlined in this paper is the result of more than five years of extensive work across various projects [16, 17, 18, 19]. It distinguishes itself from prior works in several key ways: (1) it introduces a software and hardware architecture specifically designed for deployment on a specific commercial vehicle; (2) the platform is cost-effective, making it well-suited for research and startup settings; (3) it is designed to integrate seamlessly with the vehicle’s existing systems, removing the need to add or modify the embedded systems, and thus does not require specialized engineering skills and speeding up the integration.
@@ III-A1 Vehicle type
The test vehicle is a KIA Soul EV model 2018. This vehicle has been selected because it has built-in drive-by-wire capabilities meaning that it can easily be controlled by sending actuation commands over the CAN bus. Being electric, the vehicle longitudinal control is also made easier, moreover, maintenance and operating cost are also kept low. The experimental vehicle was named Junior [17] and is shown on Fig. 1.
@@ III-A2 Vehicle interface
As the experimental vehicle is a regular commercial car, an interface providing an easy access to the CAN bus proved necessary. The DriveKit from the company Polysync [^1] https://www.polysync.io/ was used for this purpose. It is based on the Open Source Car Control (OSCC) project [20], which provides an interface to the control systems of the vehicle and allows to retrieve information such as steering wheel position or vehicle speed.
@@ III-B1 Localization
Localization techniques rely mainly on fusing different sources of information such as vehicle odometry, GNSS, IMU, and even map data [21]. Essential to get a precise vehicle localization is to find a suitable way to fuse these data to estimate the vehicle pose [22]. For simplicity and reliability, localization in Robocar is based on an all-in-one solution relying on a high-precision INS namely the Trimble BX992 capable of providing a centimeter-level positioning and a precise heading thanks to two antennas mounted on the roof. The Trimble BX992 unit estimates vehicle pose using an extended kalman filter (EKF) fusing positional information from RTK-corrected GNSS position with data coming from an embedded IMU. The unit, which runs at 20Hz, also supplies INS full-navigation data including acceleration and speed characteristics of the vehicle.
@@ III-B2 LiDAR
The LiDAR is a crucial component of the perception as it allows to construct a 3D scene around the vehicle using a point cloud, which is then used to detect and track various relevant objects in proximity of the vehicle [23] [24]. LiDARs are also crucial components to build maps [25] and help to refine the vehicle position, especially in GNSS-denied areas [26]. These sensors can be very expensive, as such, a compromise was found in the Ouster OS1. This LiDAR has a range of 100m, a vertical resolution of 64 channels and runs at 20Hz.
@@ III-B3 Cameras
Cameras complement the LiDAR perception and are also used for object detection. The experimental vehicle was equipped with six Sekonix GSML cameras providing a resolution of up to 1928x1208 at 30fps and horizontal FOVs of 60, 120 and 190 degrees. The cameras were mounted on the roof covering the front, rear and both sides of the car. An Intel RealSense D435 mounted on the windshield inside the cabin was also added and has an RGB resolution of 1920×1080 at 30fps and a horizontal FOV of 69 degrees in addition to a depth resolution up to 1280×720 at 90fps.
@@ III-B4 Onboard Computing
Being specifically designed for such purposes, the computer running the ADS is a Sintrones ABOX-5200G4 equipped with an Intel Core i7-8700T, a GTX 1060 and 32GB of RAM. This unit provides a good balance of power efficiency and cooling while remaining powerful enough for resource heavy software and moderate AI duties.
@@ III-B5 Connectivity
A TP-Link Archer MR600 router is used to provide the experimental vehicle with internet access, but also to build an internal network allowing the different computers and sensors to communicate together.
Figure 2: Summary of the cost-efficient hardware platform. The LiDAR provides a 360 degrees point cloud representation around the vehicle at a range of 100m while the Sekonix and RealSense cameras provide 120 and 69 degrees front views respectively. The onboard GNSS-INS has centimeter level accuracy.
@@ IV-A Architecture
Similar to Apollo [6], RoboCar is based on a modular architecture with modules built as ROS2 packages and containing components. A component is an individual software unit responsible for a specific task and each component is only aware of itself, thus relying on the middleware to exchange messages. A component can also periodically execute a function, using its own thread. RoboCar is multi-threaded but not multi-process, meaning that all messages are exchanged in an intra-process manner and the components constitute the ROS2 nodes. This increases the speed of message transfer, especially for larger messages such as images and point clouds. The modules can be grouped into three main categories, namely (1) the human-machine interface and services, (2) the core modules and (3) the hardware interface comprising of the drivers. Those modules are built upon the middleware, itself based on ROS2. A summary of the RoboCar software is shown in Fig. 3. The following sections will describe the different modules and their components in more detail.
@@ IV-B Middleware
The middleware is the core library used to define the components and provide them with a configuration manager allowing to define parameters and to enable or disable components in a flexible manner. The library is header-only and easy to integrate. The middleware module was initially developed as part of the FastCycle project [27] which aimed at providing a high-performance intra-process middleware. Although FastCycle was faster than ROS1, the performance were not significantly better than ROS2, as a consequence we added support for ROS2 as it provides greater flexibility and is used by a large community.
@@ IV-D Core Modules
The core modules comprises six modules essential for autonomous driving operation that do not rely on a specific hardware platform. Those modules are:
@@ IV-D1 Vehicle
This module acts as an interface between RoboCar and the experimental vehicle. A commander component continuously checks the validity of sensor data such as timestamp and accuracy, as well as other safety parameters such as steering speed, and decides accordingly of autonomous driving engagement or disengagement. If the data coming from sensors is deemed unreliable or if the vehicle is performing maneuvers outside safety limits, the system may issue warnings or completely disengage the autonomous driving mode. The commander will also send actuation commands coming from the control after a series of checks and processing.
@@ IV-D2 Localization
The localization is an essential part of the autonomous driving software as it provides the vehicle with an accurate position and orientation within a reference frame upon which the remaining core components in the autonomous driving stack rely on to function properly. The localization module converts raw GNSS positioning into an Euclidean space coordinates using a given reference, in this case, the high-definition (HD) map. In Robocar we use the Earth-North-Up (ENU) convention to get these Euclidean coordinates to represent the vehicle position with respect to the reference point. The localization may be complemented by other techniques to improve reliability in situations where the GNSS might prove insufficient [25, 21, 22, 26] or even dispense of a high precision GNSS altogether. In the scope of RoboCar and given the quality of the onboard GNSS, the localization depends solely on the output of the Trimble.
@@ IV-D3 Map
This module is responsible for parsing the different layers of an HD map and makes it available to the other core AD modules. An HD map contains lane-level geometric, semantic, and topological information of the road environment, positioned with a few centimeters of precision. [25]. It is composed of a base layer representing the 3D road environment in the format of a compressed point cloud file (.las/.laz). The geometric, semantic and topological (vector) information of the environment are represented in standard GeoJSON formats [28]. The map module is also in charge of the routing which calculates the shortest path between the current vehicle position and the destination point. To do so, the geometric and topological information of the map are used to build graph data structures so that calculating the shortest path amounts to a graph search problem for which we use the $A^{*}$ algorithm. The shortest path is then sent to the motion planning module to calculate a collision-free feasible local trajectory. Detailed implementations and other algorithms of the routing component are to be provided in future releases of Robocar.
@@ IV-D4 Perception
The perception module contains all components responsible for handling the data coming from perception sensors such as the LiDAR and the cameras. The perception is essential to provide the vehicle with a comprehensive representation of the dynamic environment surrounding it. As such, it can be tasked with detecting objects, lanes, traffic signs, traffic lights, or any other relevant features. In RoboCar, the perception module handles point clouds coming from the LiDAR as well as images coming from the cameras. The LiDAR perception component currently relies on a clustering algorithm, complemented by effective object filtering, and provides the motion planner with a set of localized objects around the vehicle. The camera perception is based on YOLOv8 [29] and handles traffic lights detection. At the time of writing, the perception system does not perform object tracking. However, efforts are in progress to implement a state-of-the-art 3D object perception pipeline based on LiDAR and camera data fusion [30], which will also include tracking capabilities.
@@ IV-D5 Planning
The motion planner is tasked with providing the control with a safe and feasible trajectory to follow under various static constraints imposed by the environment and the dynamic constraints imposed by the dynamic objects surrounding the vehicle [31, 32]. Commonly, this problem is solved using either of two approaches. In optimization-based methods, motion planning is formulated as an optimization problem, where the objective is to find the best trajectory satisfying both static and dynamic constraints. On the other hand, the sample-based motion planning methods mainly rely on randomly sampling the configuration space of the vehicle and building a graph or tree structure to explore possible paths [33]. In RoboCar, the planning module implements a novel approach based on a model predictive path integral (MPPI) algorithm. This MPPI approach uses a sample-based motion planner which generates a great number of trajectories (around 2500) and weighs them using a cost function before merging them to produce a final trajectory as depicted on Fig. 4. The work in the planning module let to [34] which thoroughly explores the implementation of the MPPI planner and evaluates it under real conditions on a closed track. Safe planning also implies respecting a safe distance at all times with other vehicles and objects which are discretized into circles as shown in green on 4. In RoboCar, the safe distance procurement is based on the responsibility-sensitive safety (RSS) model developed by Mobileye [35] in an attempt to formalize autonomous cars safety. The minimum safe distance $d_{min}$ is computed as follow :
$d_{min}=\left[d_{0}+v_{r}\rho+\frac{1}{2}a_{max}\rho^{2}+\frac{(v_{r}+\rho a_{max})^{2}}{2\beta_{min}}-\frac{v_{f}^{2}}{2\beta_{max}}\right]_{+}$ (1)
Where $d_{0}$ is an initial offset distance, $\rho$ is the reaction time, $v_{r}$ is the speed of the rear (ego) vehicle and $v_{f}$ is the speed of the front vehicle. $a_{max}$ and $\beta_{min}$ are the maximum acceleration and minimum braking of the rear vehicle while $\beta_{max}$ is the maximum braking of the front vehicle. If the actual distance between the front vehicle and the rear vehicle is less than $d_{min}$, the ego vehicle should apply some brake to restore the safe distance.
@@ IV-D6 Control
$\delta=\arctan\left(\frac{2L\sin(\alpha)}{l^{t}_{d}}\right)$ (2)
where $L$ is the wheel base and $\alpha$ the heading difference between the target point vector and the vehicle heading. The target point is located on the trajectory at a look-ahead distance $l^{t}_{d}$ computed dynamically as follow :
$l^{t}_{d}=\gamma v_{t}(1-\kappa_{t})$ (3)
Where $\gamma$ is a fixed parameter, $v_{t}$ is the current vehicle speed and $\kappa_{t}\in[0,1]$ is a normalized measurement of the curvature of the trajectory. The computed steering angle is then directly sent (after steering ratio conversion) to the actuator as the DriveKit module supports direct steering angle commands. Let $v^{r}_{t}$ and $a^{r}_{t}$ be the reference speed and acceleration respectively, and let $v^{e}_{t}=v^{r}_{t}-v_{t}$ be the measured speed error. The longitudinal control is implemented as follow :

### v2

@@ I Introduction
This research introduces RoboCar, an open-source research platform developed at the University of Luxembourg. RoboCar provides a modular, cost-effective autonomous driving solution tailored for, but not limited to, the 2018 KIA Soul EV (see Fig. 1), enabling advanced research and practical applications in autonomous driving technologies.
@@ II Related work
Similarly, the European project PRYSTINE [21] used the same vehicle platform as RoboCar to enhance automotive architectures for automated driving by focusing on fail-operational systems. As the main focus of this project was to investigate fail-operational perception based on sensor fusion, it did not implement a full autonomous driving stack, nor has the implementation been open-sourced.
Finally, the AnnieWAY team, composed of researchers from the Karlsruhe Institute of Technology and the FZI Research Center for Information Technology in Germany, participated in the 2016 Grand Cooperative Driving Challenge and developed an automated vehicle based on a modified Mercedes-Benz S-Class platform, dubbed BerthaOne. They specified a hardware and software stack for perception, cooperation, and motion planning [30]. Unfortunately, the software stack has not been open-sourced.
@@
The research outlined in this paper is the result of more than five years of extensive work across various projects [34, 35, 31, 32]. It distinguishes itself from prior works in several key ways: (1) it introduces a software and hardware architecture specifically designed for deployment on a specific commercial vehicle; (2) the platform is cost-effective, making it well-suited for research and startup settings; (3) it is designed to integrate seamlessly with the vehicle’s existing systems, removing the need to add or modify the embedded systems, and thus does not require specialized engineering skills as to speed up the integration.
@@ III Hardware Platform
Figure 2: Summary of the cost-efficient hardware platform. The LiDAR provides a 360 degrees point cloud representation around the vehicle at a range of 100m while the cameras cover all viewpoints around the vehicle. The onboard GNSS-INS has centimeter level accuracy.
@@ III-A1 Vehicle type
The test vehicle is a KIA Soul EV model 2018. This vehicle has been selected because it has built-in drive-by-wire capabilities, meaning that it can easily be controlled by sending actuation commands over the CAN bus. Being electric, the vehicle longitudinal control is also made easier, moreover, maintenance and operating cost are also kept low. The experimental vehicle was named Junior [35] and is shown on Fig. 1.
@@ III-A2 Vehicle interface
As the experimental vehicle is a regular commercial car, an interface providing an easy access to the CAN bus proved necessary. The DriveKit from the company Polysync [^1] https://www.polysync.io was used for this purpose. It is based on the Open Source Car Control (OSCC) project [24], which provides an interface to the control systems of the vehicle and allows to retrieve information such as steering wheel position or vehicle speed.
@@ III-B1 Localization
Localization techniques rely mainly on fusing different sources of information such as vehicle odometry, GNSS, IMU, and even map data [5]. Essential to get a precise vehicle localization is to find a suitable way to fuse these data to estimate the vehicle pose [29]. For simplicity and reliability, localization in Robocar is based on an all-in-one solution relying on a high-precision INS namely the Trimble BX992 capable of providing a centimeter-level positioning and a precise heading thanks to two antennas mounted on the roof. The Trimble BX992-INS unit estimates vehicle pose using an extended kalman filter (EKF) fusing positional information from RTK-corrected GNSS position with data coming from an embedded IMU. The unit, which runs at 20Hz, also supplies INS full-navigation data including acceleration and speed characteristics of the vehicle.
@@ III-B2 LiDAR
The LiDAR is a crucial component of the perception as it allows to construct a 3D scene around the vehicle using a point cloud, which is then used to detect and track various relevant objects in proximity of the vehicle [17] [7]. LiDARs are also crucial components to build maps [11] and help to refine the vehicle position, especially in GNSS-denied areas [20]. These sensors can be very expensive, as such, a compromise was found in the Ouster OS1 which provides decent performance at a fair price. This 20Hz LiDAR has a 360 degrees field of view with a range of 100m and a vertical resolution of 64 channels.
@@ III-B3 Cameras
Cameras complement the LiDAR perception and are also used for object detection. The experimental vehicle was equipped with six Sekonix GSML cameras providing a resolution of up to 1928x1208 at 30fps and horizontal FOVs of 60, 120 and 190 degrees. The cameras were mounted on the roof covering the front, rear and both sides, the purpose of this setup being to cover all viewpoints around the vehicle but it can prove to be expensive. An Intel RealSense D435 mounted on the windshield inside the cabin was also added and has an RGB resolution of 1920×1080 at 30fps and a horizontal FOV of 69 degrees in addition to a depth resolution up to 1280×720 at 90fps. The RealSense provides an alternative configuration for the camera perception where it relies on a single camera system similar to comma.ai devices [6][8]. This setup is simpler and more cost-effective but potentially limiting depending on the application. For more complex perception scenarios, a wider-angle camera than the RealSense would be best suited. If used in conjunction with the Sekonix cameras, the RealSense camera can be used instead for in-cabin monitoring.
@@ III-B4 Onboard computing
Being specifically designed for such purposes, the computer running the ADS is a Sintrones ABOX-5200G4 equipped with an Intel Core i7-8700T, a GTX 1060 and 32GB of RAM. This unit provides a good balance of power efficiency and cooling while remaining powerful enough for resource intensive software and moderate AI duties.
@@ III-B5 Connectivity
A TP-Link Archer MR600 router is used to provide the experimental vehicle with internet access, but also to build an internal network that allows the different computers and sensors to communicate together.
@@ IV-A Architecture
Similar to Apollo [3], RoboCar is based on a modular architecture with modules built as ROS2 packages and containing components. A component is an individual software unit responsible for a specific task and each component is only aware of itself, thus relying on the middleware to exchange messages. A component can also periodically execute a function, using its own thread. RoboCar is multithreaded but not multiprocess, meaning that all messages are exchanged in an intra-process manner and the components constitute the ROS2 nodes. This increases the speed of message transfer, especially for larger messages such as images and point clouds. The modules can be grouped into three main categories, namely (1) the human-machine interface and services, (2) the core modules and (3) the hardware interface comprising of the drivers. Those modules are built upon the middleware, itself based on ROS2. A summary of the RoboCar software is shown in Fig. 3. The following sections will describe the different modules and their components in more detail.
@@ IV-B Middleware
The middleware is the core library used to define the components and provide them with a configuration manager allowing to define parameters and to enable or disable components in a flexible manner. The library is header-only and easy to integrate. The middleware module was initially developed as part of the FastCycle project [31] which aimed at providing a high-performance intra-process middleware. Although FastCycle was faster than ROS1, the performance were not significantly better than ROS2, as a consequence, we added support for ROS2 as it provides greater flexibility and is used by a large community.
@@ IV-D Core modules
The core modules comprise six modules essential for autonomous driving operation that do not rely on a specific hardware platform. Those modules are:
@@ IV-D1 Vehicle
This module acts as an interface between RoboCar and the experimental vehicle. A commander component continuously checks the validity of sensor data such as timestamp and accuracy, as well as other safety parameters such as steering speed, and decides accordingly of autonomous driving engagement or disengagement. If the data coming from sensors are deemed unreliable or if the vehicle is performing maneuvers outside safety limits, the system may issue warnings or completely disengage the autonomous driving mode. The commander will also send actuation commands coming from the control after a series of checks and processing.
@@ IV-D2 Localization
The localization is an essential part of the autonomous driving software as it provides the vehicle with an accurate position and orientation within a reference frame upon which the remaining core components in the autonomous driving stack rely on to function properly. Localization in the current release of RoboCar is based on the fusion of IMU data and RTK-corrected GNSS using extended EKF embedded in the Trimble BX992-INS. This compact solution provides an accuracy of 0.05m in horizontal positioning, 0.03m in vertical positioning, and 0.09 deg. in heading in RTK mode. In GNSS denied areas, as in case of tunnels, the Trimble BX992-INS estimated pose degrades to 0.30m horizontal positioning, 0.20m in vertical positioning and 0.50 deg. in heading after 10sec of GNSS outage. While this period could tolerate short tunnels, long ones are not considered in the operational design domain (ODD) of the present release of RoboCar. To keep the safety of operation, the precision of pose estimation, e.g. position covariance is monitored and reported to anticipate takeover by the driver. Since the Trimble BX992-INS gives the results of EKF estimated pose in geodesic coordinates, the localization module converts this positioning data into Cartesian coordinates given reference point/frame, which in this case is the map reference frame. In Robocar we use the Earth-North-Up (ENU) convention to get these Cartesian coordinates to represent the vehicle position with respect to the reference point. The localization may be complemented by other techniques to improve reliability in other ODD where the GNSS might prove insufficient [11, 5, 29, 20] or even dispense of a high precision GNSS altogether. In the scope of RoboCar and given the quality of the onboard GNSS, the localization depends solely on the output of the Trimble. Using map data for localization in GNSS-denied areas is left for our future release of RoboCar.
@@ IV-D3 Map
This module is responsible for parsing the different layers of an HD map and makes it available to the other core AD modules. An HD map contains lane-level geometric, semantic, and topological information of the road environment, positioned with a few centimeters of precision. [11]. It is composed of a base layer representing the 3D road environment in the format of a compressed point cloud file (.las/.laz). The geometric, semantic and topological (vector) information of the environment are represented in standard GeoJSON formats [4]. The map module is also in charge of the routing which calculates the shortest path between the current vehicle position and a destination point. In RoboCar, the vector layers of the HD maps we provide as open source contain information on how the lane segments are connected. Each lane segment is assigned a unique identifier and the identifiers of the child and parent segments. These geometric and topological information of the map are used to build graph data structures so that calculating the shortest path amounts to a graph search problem for which we could use Dijkstra or $A^{*}$ algorithms as illustrated in Fig. 4. The shortest path is then sent to the motion planning module to calculate a collision-free feasible local trajectory. Detailed implementations and other algorithms of the routing component are to be provided in future releases of Robocar.
@@ IV-D4 Perception
The perception module contains all components responsible for handling the data coming from perception sensors such as the LiDAR and the cameras. The perception is essential to provide the vehicle with a comprehensive representation of the dynamic environment surrounding it. As such, it can be tasked with detecting objects, lanes, traffic signs, traffic lights, or any other relevant features. In RoboCar, the perception module handles point clouds coming from the LiDAR as well as images coming from the cameras. The LiDAR perception component currently relies on a clustering algorithm, complemented by effective object filtering, and provides the motion planner with a set of localized objects around the vehicle. The camera perception currently handles traffic lights detection and is based on a YOLOv8 model re-trained on the DTLD dataset [12]. The model can detect red, yellow, and green lights, and a finite-state machine in the motion planner is used to deal with potential errors and provide filtering. At the time of writing, the perception system does not perform object tracking. However, efforts are in progress to implement a state-of-the-art 3D object perception pipeline based on LiDAR and camera data fusion [19], which will also include tracking capabilities.
Figure 4: Example of shortest route calculation by the map module of RoboCar using the Belval HD map given the current vehicle position and its final destination point.
@@ IV-D5 Planning
The motion planner is tasked with providing the control with a safe and feasible trajectory to follow under various static constraints imposed by the environment and the dynamic constraints imposed by the dynamic objects surrounding the vehicle [22, 13]. Commonly, this problem is solved using either of two approaches. In optimization-based methods, motion planning is formulated as an optimization problem, where the objective is to find the best trajectory satisfying both static and dynamic constraints. On the other hand, the sample-based motion planning methods mainly rely on randomly sampling the configuration space of the vehicle and building a graph or tree structure to explore possible paths [18]. In RoboCar, the planning module implements a novel approach based on a model predictive path integral (MPPI) algorithm. This MPPI approach uses a sample-based motion planner that generates a great number of trajectories (around 2500) and weighs them using a cost function before merging them to produce a final trajectory as depicted on Fig. 5. The work in the planning module let to [32] which thoroughly explores the implementation of the MPPI planner and evaluates it under real conditions on a closed track. Safe planning also implies respecting a safe distance at all times with other vehicles and objects which are discretized into circles as shown in green on 5. In RoboCar, the safe distance procurement is based on the responsibility-sensitive safety (RSS) model developed by Mobileye [28] in an attempt to formalize autonomous cars safety. The minimum safe distance $d_{m}$ is computed as follows:
$d_{m}=\left[d_{0}+v_{r}\rho+\frac{1}{2}a_{M}\rho^{2}+\frac{(v_{r}+\rho a_{M})^{2}}{2\beta_{m}}-\frac{v_{f}^{2}}{2\beta_{M}}\right]_{+}$ (1)
where $\left[x\right]_{+}:=\text{max}(x,0)$, $d_{0}$ is an initial offset distance, $\rho$ is the reaction time, $v_{r}$ is the speed of the rear (ego) vehicle and $v_{f}$ is the speed of the front vehicle. $a_{M}$ and $\beta_{m}$ are the maximum acceleration and minimum braking of the rear vehicle while $\beta_{M}$ is the maximum braking of the front vehicle. If the actual distance between the front vehicle and the rear vehicle is less than $d_{m}$, the ego vehicle should apply some brake to restore the safe distance.
@@ IV-D6 Control
$\delta=\arctan\left(\frac{2L\sin(\alpha)}{l^{t}_{d}}\right),$ (2)
where $L$ is the wheel base and $\alpha$ the heading difference between the target point vector and the vehicle heading. The target point is located on the trajectory at a look-ahead distance $l^{t}_{d}$ computed dynamically as follows:
$l^{t}_{d}=\gamma v_{t}(1-\kappa_{t}),$ (3)
where $\gamma$ is a fixed parameter, $v_{t}$ is the current vehicle speed and $\kappa_{t}\in[0,1]$ is a normalized measurement of the curvature of the trajectory. The computed steering angle is then directly sent (after steering ratio conversion) to the actuator as the DriveKit module supports direct steering angle commands. Let $v^{r}_{t}$ and $a^{r}_{t}$ be the reference speed and acceleration respectively, and let $v^{e}_{t}=v^{r}_{t}-v_{t}$ be the measured speed error. The longitudinal control is implemented as follows: